OpenStudy (anonymous):

i am completely stuck can someone please help me. Solve. 3x2 = –21x – 15 a.7+-3sqrt69/2 b.7+-3sqrt29/2 c.-7+-sqrt29/2 d.-7+-isqrt69/2

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@sumi29

OpenStudy (whpalmer4):

Rearrange your equation so everything is on one side of the equals sign, and a 0 on the other.

OpenStudy (whpalmer4):

Having done that, see if there are any common factors to all 3 terms. If there is one, divide it out to make the rest of the problem simpler — it won't change the answer.

OpenStudy (anonymous):

ok thanks

OpenStudy (whpalmer4):

what do you get after you do that?

OpenStudy (anonymous):

well i was between c and d and i think it is d which is \[-7+i \sqrt{69}/2\]

OpenStudy (whpalmer4):

Let's work through it. What do you have for an equation after doing my suggested steps?

OpenStudy (anonymous):

hold on i threw away the paper by accident when cleaning up my work space i have to rewrite it

OpenStudy (whpalmer4):

I'll wait, no problem.

OpenStudy (anonymous):

i got \[3x ^{2}+21x+15\]

OpenStudy (whpalmer4):

=0, right? Now is there a common factor to all 3 of those terms?

OpenStudy (anonymous):

yes i forgot to add the =0 to the end

OpenStudy (whpalmer4):

common factor?

OpenStudy (anonymous):

so it would be 3 as the common factor

OpenStudy (whpalmer4):

yes. and after we take out the 3, we have?

OpenStudy (anonymous):

x^2+7x+5

OpenStudy (whpalmer4):

right. what are the values of a,b,c to put into the quadratic formula?

OpenStudy (anonymous):

now im am confused

OpenStudy (whpalmer4):

the quadratic formula says that the solutions to \[ax^2+bx+c=0,~a\ne0\] are given by \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (whpalmer4):

are you instead solving by factoring or completing the square?

OpenStudy (whpalmer4):

I guess you can't factor here, but you could complete the square.

OpenStudy (anonymous):

ok i know the second equation you put

OpenStudy (whpalmer4):

the quadratic formula. Okay, you've got \[x^2+7x+5=0\]\[ax^2+bx+c=0\]Comparing the two, what are the needed values for \(a,b,c\) to make them equal?

OpenStudy (anonymous):

zero

OpenStudy (whpalmer4):

no...look at the two equations in my previous post. what value of \(a\) makes \(ax^2=x^2\)? what value of \(b\) makes \(bx=7x\)? what value of \(c\) makes \(c = 5\)?

OpenStudy (anonymous):

by plugging in x^2 for ax^2 7x for bx and 5 for c and then i solve

OpenStudy (anonymous):

i got c am i correct

OpenStudy (anonymous):

@Nurali

OpenStudy (whpalmer4):

Yes, C is the correct choice. the solutions are \[x = \frac{1}{2}(-7\pm\sqrt{29})\]

OpenStudy (anonymous):

thanks for walking me thru this problem and being patient with me i am dumb when it comes to this stuff.